# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v24

@@ 3.19 Translators
A meta-study on the viability of machine translators replacing English to Sinhala human translators was conducted by Dilshani and Senevirathna [486]. However, this study only involves 100 combined and complex English sentences translated to Sinhala by human translators as well as MT software. Given that reason and the fact that they seem to only used Google translate and Akura Sinhala dictionary app for comparison, the conclusions of this study may not be generalized. Another meta-study on the impact of pre-trained multilingual sequence-to-sequence models on low-resource language translation has been conducted by Lee et al. [487]; while they consider Sinhala as one of the examples, they do not go much into the specific impact due to the general nature of the paper. The meta-analysis by Ramadasa et al. [488] attempts to evaluate the goodness of the Google Sinhala-to-English translation by using the Google cloud API to translate Sinhala to English and then analysing the accuracy of the Sentiment Analysis task and the Named Entity Recognition task on the translated text. The meta-analysis by Das et al. [489] compared the results of translating English to 15 Indic languages including Sinhala using statistical translation methods. They used datasets from OPUS [130] for model building and utilized Flores-200 for fine-tuning. The NMT for Indic languages study conducted by Sheshadri et al. [490] discusses the Sinhala language translation in the abstract and conclusion but the paper itself focuses more on the languages spoken in India. Nevertheless, it puts Sinhala into a regional linguistic perspective. The meta-study by Bapna et al. [491] discusses the task of building clean, web-mined datasets for a number of languages including Sinhala for the task of machine translation. This discussion was continued by Jones et al. [492] who discussed the bilingual lexica (BILEXs) in the context of a number of languages including Sinhala. Yet another meta-analysis by Halpege [493] compared the commercially available English-Sinhala translation systems provided by Google and Bing through a comprehensive error analysis. This study used academic articles as its domain and ultimately concluded that both translation systems need significant improvements before they can be recommended for widespread usage.
@@
As mentioned in Section 3.1, the study by [102] raised questions on the quality of the existing Sinhala-English corpora. In a follow-up study by Ranathunga et al. [158], it was pointed out that using an automated ranking based on a similarity measure on web-mined corpora and using the resultant top samples can yield better translation models for English-Sinhala and Sinhala-Tamil. These models sometimes were better performing than a model trained on the full dataset. They further showed that using human labour to clean web-mined parallel corpora only gives marginal benefits over automated ranking and filtering. Thus they concluded that using expensive human labour for this task might not be efficient. The study by Team et al. [494] conducted a study on the quality of the translator models trained on their NLLB data set [134]. Naturally, this includes Sinhala.
Sen et al. [495] on the other hand attempted to improve the quality of the Sinhala-English parallel corpora using fuzzy string matching where they tried to match the English translation of the given Sinhala sentence to the English sentence in the dataset pair.
@@ 3.19.2 Sinhala-English NMT
Fonseka et al. [509] used Byte Pair Encoding (BPE) for English to Sinhala neural machine translation. As another solution to the OOV problem, an analysis of subword techniques to improve English to Sinhala Neural Machine Translation (NMT) was conducted by Naranpanawa et al. [510]. A data augmentation method to expand bilingual lexicon terms based on case markers for the purpose of solving the OOV problem in the domain of NMT was proposed by Fernando et al. [511] which they later extended further [512]Epaliyana et al. [513] proposed iterative filtering and data selection be used to improve Sinhala-English NMT. Perera et al. [514] used English Part-of-Speech (PoS) tags to improve English to Sinhala NMT. Lin et al. [515] used a model based on fairseq [516] to improve machine translation between English and Sinhala. The second half of the study by Arafath [387] dealt with translating Sinhala speech to other languages. Kugathasan and Sumathipala [517] proposed an NMT system for Sinhala-English Code-Mixed text using the standardized Sinhala Code-Mixed text they proposed earlier [518]. Nguyen et al. [519] introduced a new fine-tuning objective LAgSwAV (Language-Agnostic Constraint for SwAV loss), using which they obtained 5.4 BLEU for English-Sinhala. Later, this method was further discussed by Nguyen [520]. Attigala and Weerasinghe [521] conducted an analysis on the effectiveness of ChatGPT [522] in translating Sinhala songs to English. Utsa et al. [523] used transfer learning and back translation as well as focal loss of the Sinhala-English dataset from FLoRes [98] and reported better results than mBART [524]. The study by Ranathungaa et al. [525] presents an evaluation of using parallel data from auxiliary domains to enhance EnSI NMT by fine-tuning or further pre-training the models. In their study on improving cross-lingual representation of multilingual language models for low-resource languages using linguistic entity masking, Fernando and Ranathunga [288] used SiTa and EnSi NMT on NLLB and CCAligned as one of the evaluations.
Sinhala is one of the languages considered in the short survey by Perera and Sumanathilaka [526] on machine translation and transliteration for Indo-Aryan Languages. They mention a sub-set of works [505, 509, 527, 508, 528, 529] discussed in this survey for EnSi translation.
@@ 3.19.3 Singlish to Sinhala (Transliteration)
The XTREME-UP142142142https://github.com/google-research/xtreme-up data set created by Ruder et al. [144] contains a Sinhala data set for the transliteration task. The early work by Goonetilleke et al. [530] attempted Sinhala transliteration through the Latin alphabet. However, their work does not use the word transliteration and instead focuses on the predictive aspect. Priyadarshani et al. [531] used statistical machine learning for transliteration of names between Sinhala, Tamil, and English. A rule-based system on trigrams was proposed by Liwera and Ranathunga [532] for Singlish to Sinhala transliteration of social media text. A Singlish to Sinhala converter which uses an LSTM was proposed by De Silva [533]. A rule-based approach for the same was proposed by de Silva and Ahangama [534]. The study by Nanayakkara et al. [535] introduced an English-to-Sinhala transliteration system.
@@
The work by Kirov et al. [547] used the Dakshina [157] dataset and compared the accuracy of Singlish to Sinhala transliteration with vanilla LSTM, vanilla Transformers, mT5 [336], ByT5 [548], and a non-neural finite-state transducer (FST) based on work by Bisani and Ney [549]. They discuss in detail how the zero-width joiner (ZWJ, U+200D) character cause issues in this task and Sinhala rendering as a whole. Their code is available on GitHub143143143https://github.com/google-research/google-research/tree/master/context_aware_transliteration. Khiu et al. [550] fine-tuned mBART on Sinhala Government Corpus and Bible corpus to predict the NMT performance dependence on domain similarity. De Mel et al. [551] conducted a comparative analysis for Sinhala transliteration between an extension of the rule-based approach proposed by Tennage et al. [552] and the transliteration as a translation model proposed by Deselaers et al. [553]. They reported across-the-board better performance with the latter. All thier code as well as the test-sets are publicly avaialble144144144https://github.com/kasunw22/Sinhala-Transliterator/. Perera et al. [554] attempted the same task with BERT. Their code is also publicly available145145145https://github.com/Sameera2001Perera/Singlish-Transliterator.
@@ 3.20 Spelling and Grammar
The open-source data-driven approach proposed by Wasala et al. [586, 587] claims to be able to check and correct spelling errors in Sinhala. The approach by Jayalatharachchi et al. [588] attempts to obtain synergy between two algorithms for the same purpose. These efforts [586, 588] were then extended by Subhagya et al. [589]. A rule-based Sinhala spell checker named SinSpell based on Hunspell146146146http://hunspell.github.io/ was introduced by Liyanapathirana et al. [590]. They have also made the tool available147147147http://nlp-tools.uom.lk/sinspell/ online for use. The study by Sithamparanathan and Uthayasanker [591] extended the Generic Environment for context-aware spell correction to handle Sinhala and Tamil. Sonnadara et al. [127] created a benchmark data set for Sinhala spell correction along with a neural model. A multi agent-based spell checker, named LaSi Spell for Sinhala spell checking was introduced by Samarawickrama et al. [592]. The study by Udagedara et al. [593] specifically solved the problem of spell-checking Sri Lankan names and addresses. A system named Erroff was proposed by Sudesh et al. [594] to correct real-word errors in Sinhala text. Praveen et al. [362] claim to have developed a Sinhala Spell checker with the Random Forest algorithm that archives 82% accuracy.
@@ 3.21 Chat Bots
A simple Sinhala chat bot which utilizes a small knowledge base has been proposed by Hettige and Karunananda [78]. A study on the effect of word embeddings on a Sinhala chatbot was conducted by Gamage et al. [604] where they used, the fasttext model trained by Facebook [120, 121, 122], on a RASA148148148https://rasa.com/ [605] chat bot. A Sinhala chat bot for train information was proposed by Harshani [606]. Similarly, the tool proposed by Chandrasena et al. [607] serves as a chat bot-based recommendation system for Sri Lankan traditional dancers. The chat bot discussed by Kumanayake [608] has the very specific purpose of answering user inquiries about the degree programs at University of Ruhuna. Avishka et al. [609] used off the shelf RASA NLU Engine [605] and Microsoft Bot Network [610] to set up a generic Sinhala chat bot architecture. They demonstrated the effectiveness of their architecture by creating a food ordering chat bot. A web-based code-less chat bot development platform for Sinhala was proposed by Dissanayake et al. [611]. Further, they claimed that their system can handle OOV tokens as well as Sinhala-English code-switching. The work by Dasanayaka and Warnajith [612] used a deep learning Intent Mapping (IM) model to map consumer responses in their Sinhala chat bot framework. Rajapakshe et al. [405] proposed a Sinhala conversational interface for the purpose of scheduling medical appointments and giving medical advice. The chat bot component was in the back end while the front end contained speech recognition and voice synthesizer components.
@@ 3.23 Sinhala Sign Language
Meyler [619] claim that there is no such thing as a Sinhala Sign Language or a Tamil Sign Language and there is only Sri Lankan Sign Language. Further, [620] claim that SSL was derived from British Sign Language (BSL). However, the literature uses Sri Lankan Sign Language and Sinhala Sign Language interchangeably with the same acronym, SSL. In this section, we will discuss works that have used either term. But when relevant we will point out the works that claimed to work on Sri Lankan Sign Language as opposed to Sinhala Sign Language. As far as we can observe, there does not seem to be a technical or functional difference between the two. The work by Chithrani et al. [621] compares a few Sinhala sign language recognition works [622, 623, 624, 625, 626, 627, 620, 628] against work conducted for other languages.
@@
In the domain of Sinhala sign language, Liyanaarachchi et al. [629] claims to have created a signing dataset for the Sinhala sign language however only the abstract of their work can be publicly accessible. DISSANAYAKE et al. [630] also claims to have created a database of gestures that are included in Sinhala sign language. Charuka et al. [154] has published a dynamic word-level Sinhala Sign Language video dataset of 50 classes. The dataset contains 1110 videos.
@@
Priyankara et al. [631] have conducted a short survey on the available Sinhala sign language recognition systems. They look at 12 studies and conclude that image processing with CNN is the most used technique for Sign Language translation. Along with a Sinhala sign language recognition system. Wijegoonaratna [632] has created a neural network-based approach for real-time Sinhala sign language gesture recognition. The approach by Hettiarachchi and Meegama [628], the approach by Dilakshan and Priyadarshana [633], as well as the approach by Peiris [626] have used CNN to recognize the fingerspelling alphabet of Sinhala sign language. Perera and Jayalal [634] also has used CNN to translate Sinhala sign Language to Sinhala text. However, they have also explicitly used Scale Invariant Feature Transform (SIFT) [635] features. A study for the limited use case of translating 15 Sinhala signs to text was conducted by Fernando and Wimalaratne [622]. The study by Charuka et al. [154] introduces a novel skeleton-based sign language recognition method named ProtoSign built upon prototypical networks [636] to do few-shot learning for Sinhala sign language recognition. The study by Haputhanthri et al. [637] used ResNet followed by a transformer encoder with multi-head attention to recognize sentence-level Sinhala sign language. The study by Fernando et al. [638] introduced a mobile app capable of translating Sinhala Sign language to text. They further enhance the results of their system with a facial emotion detection system created with CNNs. The study by Gamage et al. [639] also combines Sinhala Sign language recognition and emotion recognition. Krishnananthan et al. [640] introduced HANDTALK web-based application designed to teach sign-language. As Such it has a component that translates the sign language gestures of the learner into text. In their study Gedaragoda et al. [641] used traditional machine learning approaches to recognize Sinhala Sign Language. A pose-based method for Sinhala Sign Language recognition was introduced by Indatissa [642]. The work by Perera [643] uses CNN for Sri Lanka Sign Language recognition. The work by Nadeesha and Wasalthilaka [644] used classical machine learning techniques to recognize 40 Sinhala signs from 100 pose videos. They reported a success rate of 86.7%. Strides have been made in the domains of computer interpreting for written Sinhala [645] and animation of finger-spelled words and number signs [646]. The work by Idushan et al. [647] claims to be able to convert Sinhala audio to SSL and also recognize SSL and interpret dynamic signs.
@@
The work by Samarasinghe et al. [623] used a CNN model to convert Sinhala sign language to Sinhala text. Kumar et al. [648] introduced EasyTalk a tool that translates Sri Lankan sign language to text as well as converts English text into Sri Lankan Sign Language. The work by dha [627] used CNN to classify features of static Sinhala signs while dynamic signs were modelled as a tree structure of static signs. Dahanayaka et al. [625] also used CNN in their multi-model approach to recognize Sinhala sign language. Both the approaches by Perera and Kumarika [649] and Herath and Ishanka [650] integrated MediaPipe [651] with Deep Neural Networks to interpret Sri Lankan sign language. Perera and Kumarika [649] only used LSTM while Herath and Ishanka [650] used LSTM, CNN, and CNN-LSTM. Rishan et al. [620] uses Leap Motion Controller149149149https://leap2.ultraleap.com/leap-motion-controller-2/ to translate Sri Lankan sign language to Sinhala text. Dissanayake et al. [624] have introduced Utalk, a mobile app to interpret Sri Lankan Sign Language.

### v25

@@ 3.19 Translators
A meta-study on the viability of machine translators replacing English to Sinhala human translators was conducted by Dilshani and Senevirathna [502]. However, this study only involves 100 combined and complex English sentences translated to Sinhala by human translators as well as MT software. Given that reason and the fact that they seem to only used Google translate and Akura Sinhala dictionary app for comparison, the conclusions of this study may not be generalized. Another meta-study on the impact of pre-trained multilingual sequence-to-sequence models on low-resource language translation has been conducted by Lee et al. [503]; while they consider Sinhala as one of the examples, they do not go much into the specific impact due to the general nature of the paper. The meta-analysis by Ramadasa et al. [504] attempts to evaluate the goodness of the Google Sinhala-to-English translation by using the Google cloud API to translate Sinhala to English and then analysing the accuracy of the Sentiment Analysis task and the Named Entity Recognition task on the translated text. The meta-analysis by Das et al. [505] compared the results of translating English to 15 Indic languages including Sinhala using statistical translation methods. They used datasets from OPUS [130] for model building and utilized Flores-200 for fine-tuning. The NMT for Indic languages study conducted by Sheshadri et al. [506] discusses the Sinhala language translation in the abstract and conclusion but the paper itself focuses more on the languages spoken in India. Nevertheless, it puts Sinhala into a regional linguistic perspective. The meta-study by Bapna et al. [507] discusses the task of building clean, web-mined datasets for a number of languages including Sinhala for the task of machine translation. This discussion was continued by Jones et al. [508] who discussed the bilingual lexica (BILEXs) in the context of a number of languages including Sinhala. Yet another meta-analysis by Halpege [509] compared the commercially available English-Sinhala translation systems provided by Google and Bing through a comprehensive error analysis. This study used academic articles as its domain and ultimately concluded that both translation systems need significant improvements before they can be recommended for widespread usage. The survey by Tafa et al. [510] compared the performance of a number of MT models on Low-Resource languages, including Sinhala.
@@
As mentioned in Section 3.1, the study by [102] raised questions on the quality of the existing Sinhala-English corpora. In a follow-up study by Ranathunga et al. [158], it was pointed out that using an automated ranking based on a similarity measure on web-mined corpora and using the resultant top samples can yield better translation models for English-Sinhala and Sinhala-Tamil. These models sometimes were better performing than a model trained on the full dataset. They further showed that using human labour to clean web-mined parallel corpora only gives marginal benefits over automated ranking and filtering. Thus they concluded that using expensive human labour for this task might not be efficient. The study by Team et al. [511] conducted a study on the quality of the translator models trained on their NLLB data set [134]. Naturally, this includes Sinhala. Sen et al. [512] on the other hand attempted to improve the quality of the Sinhala-English parallel corpora using fuzzy string matching where they tried to match the English translation of the given Sinhala sentence to the English sentence in the dataset pair.
@@ 3.19.2 Sinhala-English NMT
Fonseka et al. [526] used Byte Pair Encoding (BPE) for English to Sinhala neural machine translation. As another solution to the OOV problem, an analysis of subword techniques to improve English to Sinhala Neural Machine Translation (NMT) was conducted by Naranpanawa et al. [527]. A data augmentation method to expand bilingual lexicon terms based on case markers for the purpose of solving the OOV problem in the domain of NMT was proposed by Fernando et al. [528] which they later extended further [529]Epaliyana et al. [530] proposed iterative filtering and data selection be used to improve Sinhala-English NMT. Perera et al. [531] used English Part-of-Speech (PoS) tags to improve English to Sinhala NMT. Lin et al. [532] used a model based on fairseq [533] to improve machine translation between English and Sinhala. The second half of the study by Arafath [400] dealt with translating Sinhala speech to other languages. Kugathasan and Sumathipala [534] proposed an NMT system for Sinhala-English Code-Mixed text using the standardized Sinhala Code-Mixed text they proposed earlier [535]. Nguyen et al. [536] introduced a new fine-tuning objective LAgSwAV (Language-Agnostic Constraint for SwAV loss), using which they obtained 5.4 BLEU for English-Sinhala. Later, this method was further discussed by Nguyen [537]. Attigala and Weerasinghe [538] conducted an analysis on the effectiveness of ChatGPT [539] in translating Sinhala songs to English. Utsa et al. [540] used transfer learning and back translation as well as focal loss of the Sinhala-English dataset from FLoRes [98] and reported better results than mBART [541]. The study by Ranathungaa et al. [542] presents an evaluation of using parallel data from auxiliary domains to enhance EnSI NMT by fine-tuning or further pre-training the models. In their study on improving cross-lingual representation of multilingual language models for low-resource languages using linguistic entity masking, Fernando and Ranathunga [292] used SiTa and EnSi NMT on NLLB and CCAligned as one of the evaluations. Thillainathan et al. [543] proposed to use continual pre-training and intermediate task transfer learning for Fine-tuning what they call multilingual sequence-to-sequence large language models (msLLMs). However, in actuality, what their experiments are on is mBART.
Fernando et al. [544] in their meta-study on low-resource language NMT, introduced debiasing heuristics for web mined parallel corpora that improve EnSi translation quality. They showed the improvements on CCMatrix and CCAligned datasets. The meta study by Velayuthan et al. [545] proposes an improved sequence level Knowledge Distillation method that improves encoder knowledge transfer. They used Sinhala-English NMT to discuss the $\alpha$ value ablation by using CCAligned [546], OpenSubtitles [133], and SITA (Gov) [524] corpora. Sinhala is one of the languages considered in the short survey by Perera and Sumanathilaka [547] on machine translation and transliteration for Indo-Aryan Languages. They mention a sub-set of works [522, 526, 548, 525, 549, 550] discussed in this survey for EnSi translation.
@@ 3.19.3 Singlish to Sinhala (Transliteration)
The XTREME-UP147147147https://github.com/google-research/xtreme-up data set created by Ruder et al. [144] contains a Sinhala data set for the transliteration task. The early work by Goonetilleke et al. [551] attempted Sinhala transliteration through the Latin alphabet. However, their work does not use the word transliteration and instead focuses on the predictive aspect. Priyadarshani et al. [552] used statistical machine learning for transliteration of names between Sinhala, Tamil, and English. A rule-based system on trigrams was proposed by Liwera and Ranathunga [553] for Singlish to Sinhala transliteration of social media text. A Singlish to Sinhala converter which uses an LSTM was proposed by De Silva [554]. A rule-based approach for the same was proposed by de Silva and Ahangama [555]. The study by Nanayakkara et al. [556] introduced an English-to-Sinhala transliteration system.
@@
The work by Kirov et al. [568] used the Dakshina [157] dataset and compared the accuracy of Singlish to Sinhala transliteration with vanilla LSTM, vanilla Transformers, mT5 [349], ByT5 [569], and a non-neural finite-state transducer (FST) based on work by Bisani and Ney [570]. They discuss in detail how the zero-width joiner (ZWJ, U+200D) character cause issues in this task and Sinhala rendering as a whole. Their code is available on GitHub148148148https://github.com/google-research/google-research/tree/master/context_aware_transliteration. Khiu et al. [571] fine-tuned mBART on Sinhala Government Corpus and Bible corpus to predict the NMT performance dependence on domain similarity. De Mel et al. [572] conducted a comparative analysis for Sinhala transliteration between an extension of the rule-based approach proposed by Tennage et al. [573] and the transliteration as a translation model proposed by Deselaers et al. [574]. They reported across-the-board better performance with the latter. All thier code as well as the test-sets are publicly avaialble149149149https://github.com/kasunw22/Sinhala-Transliterator/. Perera et al. [575] attempted the same task with BERT. Their code is also publicly available150150150https://github.com/Sameera2001Perera/Singlish-Transliterator.
@@ 3.20 Spelling and Grammar
The open-source data-driven approach proposed by Wasala et al. [607, 608] claims to be able to check and correct spelling errors in Sinhala. The approach by Jayalatharachchi et al. [609] attempts to obtain synergy between two algorithms for the same purpose. These efforts [607, 609] were then extended by Subhagya et al. [610]. A rule-based Sinhala spell checker named SinSpell based on Hunspell151151151http://hunspell.github.io/ was introduced by Liyanapathirana et al. [611]. They have also made the tool available152152152http://nlp-tools.uom.lk/sinspell/ online for use. The study by Sithamparanathan and Uthayasanker [612] extended the Generic Environment for context-aware spell correction to handle Sinhala and Tamil. Sonnadara et al. [127] created a benchmark data set for Sinhala spell correction along with a neural model. A multi agent-based spell checker, named LaSi Spell for Sinhala spell checking was introduced by Samarawickrama et al. [613]. The study by Udagedara et al. [614] specifically solved the problem of spell-checking Sri Lankan names and addresses. A system named Erroff was proposed by Sudesh et al. [615] to correct real-word errors in Sinhala text. Praveen et al. [375] claim to have developed a Sinhala Spell checker with the Random Forest algorithm that archives 82% accuracy.
@@ 3.21 Chat Bots
A simple Sinhala chat bot which utilizes a small knowledge base has been proposed by Hettige and Karunananda [78]. A study on the effect of word embeddings on a Sinhala chatbot was conducted by Gamage et al. [625] where they used, the fasttext model trained by Facebook [120, 121, 122], on a RASA153153153https://rasa.com/ [626] chat bot. A Sinhala chat bot for train information was proposed by Harshani [627]. Similarly, the tool proposed by Chandrasena et al. [628] serves as a chat bot-based recommendation system for Sri Lankan traditional dancers. The chat bot discussed by Kumanayake [629] has the very specific purpose of answering user inquiries about the degree programs at University of Ruhuna. Avishka et al. [630] used off the shelf RASA NLU Engine [626] and Microsoft Bot Network [631] to set up a generic Sinhala chat bot architecture. They demonstrated the effectiveness of their architecture by creating a food ordering chat bot. A web-based code-less chat bot development platform for Sinhala was proposed by Dissanayake et al. [632]. Further, they claimed that their system can handle OOV tokens as well as Sinhala-English code-switching. The work by Dasanayaka and Warnajith [633] used a deep learning Intent Mapping (IM) model to map consumer responses in their Sinhala chat bot framework. Rajapakshe et al. [418] proposed a Sinhala conversational interface for the purpose of scheduling medical appointments and giving medical advice. The chat bot component was in the back end while the front end contained speech recognition and voice synthesizer components. The end product of work by Dasanayaka et al. [634] is a Sinhala chatbot that can answer patient questions on orthopantomography radiology reports which were previously generated by the system itself by analysing dental panoramic tomography (DPT) images.
@@ 3.23 Sinhala Sign Language
Meyler [641] claim that there is no such thing as a Sinhala Sign Language or a Tamil Sign Language and there is only Sri Lankan Sign Language. Further, Rishan et al. [642] claim that SSL was derived from British Sign Language (BSL). However, the literature uses Sri Lankan Sign Language and Sinhala Sign Language interchangeably with the same acronym, SSL. In this section, we will discuss works that have used either term. But when relevant we will point out the works that claimed to work on Sri Lankan Sign Language as opposed to Sinhala Sign Language. As far as we can observe, there does not seem to be a technical or functional difference between the two. The work by Chithrani et al. [643] compares a few Sinhala sign language recognition works [644, 645, 646, 647, 648, 649, 642, 650] against work conducted for other languages.
@@
In the domain of Sinhala sign language, Liyanaarachchi et al. [651] claims to have created a signing dataset for the Sinhala sign language however only the abstract of their work can be publicly accessible. DISSANAYAKE et al. [652] also claims to have created a database of gestures that are included in Sinhala sign language. Charuka et al. [154] has published a dynamic word-level Sinhala Sign Language video dataset of 50 classes. The dataset contains 1110 videos. Abhishek and Sumanathilaka [653] claims to introduce SSL400, a word-level Sinhala sign language dataset consisting of 384 commonly used words. However, given that their paper is behind a paywall, it is not possible to verify whether the dataset is publicly available or not.
@@
Priyankara et al. [654] have conducted a short survey on the available Sinhala sign language recognition systems. They look at 12 studies and conclude that image processing with CNN is the most used technique for Sign Language translation. Along with a Sinhala sign language recognition system. Wijegoonaratna [655] has created a neural network-based approach for real-time Sinhala sign language gesture recognition. The approach by Hettiarachchi and Meegama [650], the approach by Dilakshan and Priyadarshana [656], as well as the approach by Peiris [648] have used CNN to recognize the fingerspelling alphabet of Sinhala sign language. Perera and Jayalal [657] also has used CNN to translate Sinhala sign Language to Sinhala text. However, they have also explicitly used Scale Invariant Feature Transform (SIFT) [658] features. A study for the limited use case of translating 15 Sinhala signs to text was conducted by Fernando and Wimalaratne [644]. The study by Charuka et al. [154] introduces a novel skeleton-based sign language recognition method named ProtoSign built upon prototypical networks [659] to do few-shot learning for Sinhala sign language recognition. The study by Haputhanthri et al. [660] used ResNet followed by a transformer encoder with multi-head attention to recognise sentence-level Sinhala sign language. The study by Fernando et al. [661] introduced a mobile app capable of translating Sinhala Sign language to text. They further enhance the results of their system with a facial emotion detection system created with CNNs. The study by Gamage et al. [662] also combines Sinhala Sign language recognition and emotion recognition. Krishnananthan et al. [663] introduced HANDTALK web-based application designed to teach sign-language. As Such it has a component that translates the sign language gestures of the learner into text. In their study Gedaragoda et al. [664] used traditional machine learning approaches to recognize Sinhala Sign Language. A pose-based method for Sinhala Sign Language recognition was introduced by Indatissa [665]. The work by Perera [666] uses CNN for Sri Lanka Sign Language recognition. The work by Nadeesha and Wasalthilaka [667] used classical machine learning techniques to recognize 40 Sinhala signs from 100 pose videos. They reported a success rate of 86.7%. Rajapakshe et al. [668] used a hybrid CNN-LSTM model and reports 94% accuracy on the 20 Sinhala signs they tested on. The work by Navodya et al. [669] reports accuracies between 87% and 92% for recognising Sinhala sign language. Thennakon et al. [670] on the other hand reports 96% accuracy in recognising real-time Sinhala Sign Language signs. The work by Kavinda et al. [671] proposed a multi-modal system that converts both sign language and miming language to Sinhala text. Further, they have a component to convert Sinhala audio to miming language and Sinhala text as well. Strides have been made in the domains of computer interpreting for written Sinhala [672] and animation of finger-spelt words and number signs [673]. The survey by Ahinsa et al. [674] discusses text-to-Sinhala sign language in the contexts of: finger-spelling and translation. It should be noted here that the translation they describe here is text-to-sign language translation and not translation between any two or more natural languages. The work by Idushan et al. [675] claims to be able to convert Sinhala audio to SSL and also recognize SSL and interpret dynamic signs.
@@
The work by Samarasinghe et al. [645] used a CNN model to convert Sinhala sign language to Sinhala text. Kumar et al. [676] introduced EasyTalk, a tool that translates Sri Lankan sign language to text as well as converts English text into Sri Lankan Sign Language. The work by dha [649] used CNN to classify features of static Sinhala signs, while dynamic signs were modelled as a tree structure of static signs. Dahanayaka et al. [647] also used CNN in their multi-model approach to recognize Sinhala sign language. Both the approaches by Perera and Kumarika [677] and Herath and Ishanka [678] integrated MediaPipe [679] with Deep Neural Networks to interpret Sri Lankan sign language. Perera and Kumarika [677] only used LSTM while Herath and Ishanka [678] used LSTM, CNN, and CNN-LSTM. Rishan et al. [642] uses Leap Motion Controller154154154https://leap2.ultraleap.com/leap-motion-controller-2/ to translate Sri Lankan sign language to Sinhala text. Dissanayake et al. [646] have introduced Utalk, a mobile app to interpret Sri Lankan Sign Language.